OpenStudy (anonymous):

write 87% and 29.8% as decimals

OpenStudy (unklerhaukus):

\[87\%=\frac{87}{100}=\frac{8.7}{10}=\frac{\dots}{1}=\quad\dots\]

OpenStudy (anonymous):